a truck traveling 45 mph and a train traveling 60 mph cover the same distance . the trucks travels 2 hours longer than the train. how many does each travel
Answer by bam878s(77)

we know distance = rate x time
for the truck, we have distance = 45 x (t+2). Letting t = time.
and for the train we have distance = 60 x t.
So, we have,
45(t+2) = 60t
45t + 90 = 60t
90 = 15t
t = 6.
now plug this value into one of the original equations to solve for the distance.
distance = 60t = 60(6) = 360 miles.